A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective.

Step 2 of 2:

Suppose a sample of 715 floppy disks is drawn. Of these disks, 651 were not defective. Using the data, construct the 90% confidence interval for the population proportion of disks which are defective. Round your answers to three decimal places.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 651 / 715 = 0.910

1 - \hat p = 1 - 0.910 = 0.09

Z\alpha/2 = 1.645

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.645 * (\sqrt((0.910 * 0.09) / 715)

= 0.018

A 90% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.910 - 0.018 < p < 0.910 + 0.018

0.892 < p < 0.928

The 90% confidence interval for the population proportion p is : 0.892 , 0.928